OpenStudy (anonymous):

We roll a pair of dice. If the sum of the dice is 7, you pay me $32. If the sum is not 7, I pay you the number of dollars indicated by the sum of the dice. What is your expected value for the game?

OpenStudy (perl):

how many ways can you roll to 7?

OpenStudy (perl):

(die1, die2) 1,6 2,5 3,4 4,3 5,2 6,1

OpenStudy (perl):

EV = x1*P(x1) + x2*P(x2) + .. xn*P(xn)

OpenStudy (perl):

if i roll a sum of 7, i lose 32 dollars. if i dont roll a sum of 7 , i gain the sum of the dice

OpenStudy (perl):

if i roll a sum of 7, i lose 32 dollars. if i dont roll a sum of 7 , i gain the dollar amount of the sum of the dice

OpenStudy (anonymous):

mmkay do what @perl says XD

OpenStudy (perl):

now you have to break down the cases

OpenStudy (dan815):

i think there might be a nice symmetry im ust gonna guess before seeing that full form pan out

OpenStudy (perl):

there is no trick or fancy math required here, its purely mechanical

OpenStudy (dan815):

7= 1/6th of the time 5/6th of the time im still gonna get expected sum = 7 so 7*5=35m and u are gaining 32, so on ag u lose 3 bucks

OpenStudy (perl):

there are sums 2,3,4,5,6,8,9,10,11,12 (note there is no 7)

OpenStudy (dan815):

im thinkinngg expected is still 7, not sure though :P

OpenStudy (perl):

im going to upload an image

OpenStudy (perl):

damn, attach file button is disabled

OpenStudy (perl):

Here are the probability of P(X) , where x = sum of two die P(2)=1/36 P(3)=2/36 P(4)=3/36 P(5)=4/36 P(6)=5/36 P(7)=6/36 P(8)=5/36 P(9)=4/36 P(10)=3/36 P(11)=2/36 P(12)=1/36

OpenStudy (perl):

EV = (-7)P(7) + 2*P(2) + 3*P(3) + 4*P(4) + 5*P(5) + 6*P(6) + 8*P(8) + 9*P(9) + 10*P(10) + 11*P(11) + 12*P(12)

OpenStudy (perl):

and chucky's gone

OpenStudy (perl):

EV = 2*1/36 + 3*2/36 + 4*3/36 + 5*4/36 + 6*5/36 + (-7)*6/36 + 8*5/36 + 9*4/36 + 10*3/36 + 11*2/36 +12*1/36

OpenStudy (perl):

EV = 14/3 = 4.16666666666666666666666666666666666666666666666666

OpenStudy (dan815):

hmm

OpenStudy (dan815):

feels weird

OpenStudy (dan815):

i feel like he shud be gaining 3 bucks on avg

OpenStudy (dan815):

why did my symmetry fail

OpenStudy (perl):

4.67 dollars gain on average

OpenStudy (perl):

the dollar amounts you win are weighted by the probability of getting that sum

OpenStudy (dan815):

right look at this argument

OpenStudy (perl):

they are not uniform,

OpenStudy (dan815):

1/36 times its a 2 1/36 times in a 12 on avg 2/36 times u are paying a 12 or 2 so its avg of 7 bucks

OpenStudy (perl):

ok

OpenStudy (dan815):

similarliy 2/36 times u pay 3 2/36 times u have 11 again on avg 4/36 u pay an avg of 7

OpenStudy (dan815):

that shud continue till 8 and 6

OpenStudy (dan815):

then -7*6/36

OpenStudy (perl):

interesting :)

OpenStudy (dan815):

i mean 32*

OpenStudy (dan815):

5/6th of the time u shud be getting 7 and 1/6th u lose 32

OpenStudy (dan815):

something must be wrong with my arguments

OpenStudy (perl):

ill see what i can find on paper, sometimes things pop out on paper

OpenStudy (perl):

you dont mind if i get back to you. or if anyone else wants to try to solve dan's symmetry argument @everyone

OpenStudy (dan815):

lol

OpenStudy (perl):

iheartmath @ Dan

OpenStudy (perl):

dan, youre pretty creative. you did that without knowing about EV?

OpenStudy (perl):

you have an intuition for probability

OpenStudy (dan815):

i knew the expected value to work but just thought this was interesting lol

OpenStudy (perl):

oh ok

OpenStudy (dan815):

kinda lazy to write out all the expected values

OpenStudy (perl):

yeah it was a chore

OpenStudy (dan815):

and thanks <3

OpenStudy (perl):

i had to put it in notepad

OpenStudy (perl):

maybe i made a mistake in arithmetic?

OpenStudy (dan815):

no its probably righ

OpenStudy (perl):

this is cool question

OpenStudy (dan815):

@ganeshie8 ganeshie likes this kinda stuff too

OpenStudy (dan815):

oh you stopped going there now?

OpenStudy (dan815):

alright bye baby

OpenStudy (dan815):

im not sure if this helps in anyway to think about it like this

OpenStudy (dan815):

if u imagine a binomial distribution 2 to 12, and 7 being the mean, removing the mean still has same ammount left and right of the mean , so 7 should still be the ev

OpenStudy (dan815):

ganeshie can u tell me why my way of thinking fails

OpenStudy (dan815):

http://prntscr.com/50uz55 this part

OpenStudy (perl):

i like problems that have 'mechanical' solutions, though

OpenStudy (perl):

im not so great at the creativity part

OpenStudy (dan815):

oh i see it makes sense now!!

OpenStudy (dan815):

it shud be skewed more to the right, since the weight is shifting to the right ofc!

OpenStudy (dan815):

ugh nvm

OpenStudy (perl):

is that your final answer?

OpenStudy (dan815):

nope, we ganeshie will tell me why im wrong :)

OpenStudy (dan815):

look it does make sense if u eliminate!! the -7 every 1/6th time

OpenStudy (dan815):

out of the 30 rolls, u do get a 7 on avg

OpenStudy (dan815):

7*30/36 - 32*6/26 =[ why this fails!!

ganeshie8 (ganeshie8):

the probability for each sum is not same rihgt ?

OpenStudy (dan815):

yeah

OpenStudy (perl):

the values to the right are greater, so will make it seesaw to the right

ganeshie8 (ganeshie8):

7 has the highest probabilty because a+b=7 number of ways of choosing two positive integers for a, b = 6 since a,b can only be chosen from {1,2,3,4,5,6}, we are getting some symmetry: P(n) = P(14-n) hmm

OpenStudy (dan815):

i wonder if we can do think about it like stick and placement problem to see how much of each sum shud exist

ganeshie8 (ganeshie8):